In a physics lab, a 0.500-kg cart (Cart A) moving with a speed of 130 m/s encounters a magnetic collision with a 1.50-kg cart (Cart B) which is initially at rest. The 0.500-kg cart rebounds with a speed of 50 m/s in the opposite direction. Determine the post-collision speed of the 1.50-kg cart.
Answer #1

using conservation of momentum

m1v1i + m2v2i = m1v1f + m2v2f

0.5*130 + 0 = 0.5(-50) + 1.5*v2f

65 = - 25 + 1.5*v2f

v2f = 60 m/s